Home‎ > ‎Reports‎ > ‎Crew101‎ > ‎

day09


Crew 101 Log Book for March 20, 2011

Photos of the day





Commander's Check-In Report

Tom Caillouet Reporting

Crew Health: Quite good

EVAs and Times:

EVA 10: 1155 - 1405

EVA 11: 1643 - 1728

EVA 12: 1800 - 1915

Number of Reports: 9 (Commander, Engineering, Journalist, Chef, 3 EVA, 2 Science)

Field Results:

EVA 10 - Successfully mapped radio coverage on Cow Dung Road

EVA 11 - Checked radio repeater functionality, return to Hab with wires to fix

EVA 12 - Replaced fixed wires, raised repeater antenna

Plans for Tomorrow: Morning EVA to test radio coverage on road to radio ridge (early scouting suggests limited range). Afternoon EVA for geo/bio sample gathering.


Commander's Report

Tom Caillouet Reporting

Wind looks to be a big problem for us for the second week we are to be on Mars. It seems like we have gotten off alright in losing the wind turbine; the high winds we are expecting to get over the next few days are much higher than it has been run in the past. With this weather, we likely would have had to pull down the turbine for safety reasons. All other science has been going well. Earlier today, Cody and Chrissy rode the length of Cow Dung Road to test radio transmission strength.

The afternoon EVAs were climbs up Radio Ridge for Jon and Mitesh to check on the radio repeater and increase antenna height. Inside the Hab, Christina has been hard at work testing the biology samples with some help from Mitesh. I even spent some time playing scientist with the geology samples, figuring the percentages of the samples of different dust sizes. It has been a very productive day.


Engineering Report

Cody Hall Reporting

Generator/Electricity:

• Kitty (Diesel Generator) (Used? Yes/No): Yes

• Honey (Gas Generator) (Used? Yes/No): No

• Battery Charge Level:

o V everstart

o V armor plate

o V green

Notes/Comments (include how many generator runs, approximately when,

and any times you needed to turn the generator on/off):

2 generator runs


Fuel Status:

• Diesel (Full/0.75/0.5/0.25/Empty): 0.75

• Propane Tank (Full/0.75/0.5/0.25/Empty): 0.25

• Gasoline Tank (Full/0.75/0.5/0.25/Empty): 0.5

Notes/Comments:


Oil:

• Kitty Oil Quantity (# of Quarts in storage): Changed/replaced by DG

• Honey Oil Quantity (# of Quarts in storage): Leave it to DG

• ATV Oil Quantity (# of Quarts in storage): 6 quarts on the

engineering work bench

• Car Oil Quantity (# of Quarts in storage): Leave it to DG

Notes/Comments:


Water Status:

• Outside Potable Water Tank Level (inches from bottom): 19

• Trailer Potable Water Tank Level (inches from bottom): DG

will be returning to us tonight after refilling

• Hab Potable Water Tank Level (inches from bottom): 8

• Potable Water Meter Reading: 437026

Notes/Comments :


GreenHab:

• Condition of plants in Tank 1: unchanged

• Condition of plants in Tank 2: unchanged

• Condition of plants in Tank 3: unchanged

• Greenhab notes/ comments:


Observatory:

• Telescope Used? (Yes/no) No

Observatory Notes:


Transportation:

Hab Car (NOT RENTAL)

• Used (yes/no): yes

• Oil Checked (yes/no): No

• Tire Status: Not checked

• Odometer at end of day: 123688.4

• Notes/Comments on Hab Car: ran for 5 minutes on recommendation

of DG to check status and charge battery. started fine.


ATV 1 (Opportunity):

• Used: yes

• Oil Checked: no

• Fuel Consumed: replaced

• Tire Status: ok

• Comments:


ATV 2 (Spirit):

• Used: yes

• Oil Checked: no

• Fuel Consumed: replaced

• Tire Status: ok

• Comments:


ATV 3 (Viking I):

• Used: no

• Oil Checked: no

• Fuel Consumed: none

• Tire Status: ok

• Comments


ATV 4 (Viking II):

• Used: no

• Oil Checked: no

• Fuel Consumed: none

• Tire Status: ok

• Comments:


ATV Comments:


Heating and Ventilation:

• Thermostat setting upstairs (Farenheit): 68

• Actual temperature upstairs (Farenheit): 69

• Thermostat setting downstairs (Farenheit): 65

• Actual temperature downstairs (Farenheit): 65

Notes/Comments:


Computers/Networking Infrastructure: working

Notes/comments on internet/computers:


General Engineering

Notes/Comments:


Science Report

Tom Caillouet Reporting

After a less than completely successful EVA to measure radio signal the radio repeater antenna. The morning EVA was, however, successful, wind appeared to have minimal impact on the results.

strength on March 18th, plans were made to once again map signal strength along Cow Dung Road on the 20th. There was concern due to increased wind activity; this could lead to reduced clarity by causing interference in EVA crews' microphonse or reduced power by deforming

 The procedure was to send ATVs along the trail and pause every couple minutes to report in GPS coordinates and altitude. A radio operator at the Hab would record this data, along with a signal strength and clarity rating between 1 and 5, with 1 being no strength or clarity and 5 being perfect strength or clarity. If the Hab could not be reached from the road, the EVA crew would climb a nearby hill to attempt to boost the signal. Ratings for the route follow, with site 1 being a reading at the base of the Hab:


Site             UTM Coordinates                Altitude (m)   Strength         Clarity        On Road?

1               12N 0518229 4250721             1372            5                       5                       Y

2               12N 0518320 4250573             1364            5                       5                       Y

3               12N 0518873 4250922             1368            5                       4                       Y

4               12N 0519180 4251734             1370            5                       4                       Y

5               12N 0518830 4252567             1372            5                       4                       Y

6               12N 0518561 4253033             1369            5                       4                       N

7               12N 0517769 4254467             1383            5                       3                       N

8               12N 0518065 4255477             1389            5                       2                       N

9               12N     0517762 4254866         1387            5                       3                       N

10              12N 0518054 4253827             1373            5                       3                       N

11              12N 0518230 4250097             1360            5                       5                       Y

12              12N 0518485 4249102             1354            5                       4                       Y

13              12N 0519052 4248400             1352            5                       3                       N

14              12N 0519903 4247996             1347            5                       3                       N

15              12N 0520345 4247080             1344            5                       3                       Y


The first obvious trend is that the signal strength is perfect for all transmissions. This is due to the radio repeater located above the Hab. Any loss in signal power is lost between the EVA crew and the repeater, then the repeater boosts the signal to transmit to the Hab. Therefore, the radio operator at the Hab will never detect a signal power fluctuation, the signal strength will either appear to be 100% or completely lost.


Clarity, on the other hand varies rather significantly. There is no apparent loss in clarity within 500 - 700 meters of the Hab. Up to about three kilometers north of the Hab, there is minimal noise in the transmission. Southern bound crews start to gain additional noise around 2.5 kilometers south of the Hab, likely due to interference caused by Kissing Camel Range. Full communications range is around 5 kilometers from the Hab, though crews often had to climb nearby hills to gain two-way communications. This is partially because the northbound trail often has small hills just off the path blocking transmission. These seem to have a greater effect on blocking signal than larger features that are further removed from the trail, such as Olympus Mons.

This 5 kilometer radius of transmissions theoretically gives EVA crews a 157 square kilometers of imperfect radio coverage. Geographic features do clearly disrupt the signals, but positioning the radio equipment in front of or on top of hills allows for contact with the Hab. Radio signal on top of Radio Ridge, which accounts for about half of that area, is also untested at the moment. We hope to test signal strength in that area tomorrow, if weather conditions permit.


Science Report (Biology)

Christina Graves Reporting

Objectives:

Science objectives for the day is to make all solutions necessary for E.Z.N.A. Soil DNA (Omega) kit, and to gather all equipment necessary for the E.Z.N.A. Soil DNA Protocol, as well as for the chromatography experiments. In addition, soil sample solutions will be plated on premade, sterile Agar plates (Sigma-Aldrich).


Protocol:

First, 10-2 dilutions were made of all samples collected thus far. Suggested dilution was 5g soil to 450mL of dIH2O.

For 15mL of dIH2O, used 167mg of soil. (balance sensitivity – 0.18g)

Used Corning 15mL conical tubes.

031911 – 3 samples

1) 51.40g (referred to as Sample 7)

2) 222.66g – dilution not made

3) 139.69g (referred to as Sample 8)


031811 – 3 samples

1) 204.25g (referred to as Sample 5)

2) 142.29g (referred to as Sample 6)

3) large rock – dilution not made


031711 – 1 sample

1) 320.8g (referred to as Sample 4)


031511 – 3 samples

1) 108.33g (referred to as Sample 1)

2) 220.98g (referred to as Sample 2)

3) 463.64g (referred to as Sample 3)


2. Make 10-3 and 10-4 dilutions of stocks, using 20uL of 10-2 and 10-3 solutions (respectively) to 180uL of dIH2O.


3. Prepare adenine, cytosine, and thymine base stock solutions for chromatography experiments.

200mg each base to 150mL dIH2O.


These have different solubilities, but concentration of the standards is not important or this experiment , as well will be performing qualitative and not quantitative comparisons for the experiment.


Note: guanine not prepared because it has extremely low solubility in water.

Heat solutions over hot water to increase dissolution of base into water.

Make dilutions until clear liquid remains.

Add 20uL of super-saturated solution to 180uL of dIH2O.


3. Calibrate thermocouple for use for the E.Z.N.A. Soil kit, to be performed tomorrow.

4. Make ice for ice bath (freeze overnight).

5. Put water in water bath - this does not seem to be working - will allow to heat up for a few hours. May need to use hot water.

6. Turn on incubators to prepare for experiment tomorrow. Set at 70C.

7. Plate 100uL of each 10-3 sample (labelled Samples 1-8) on a premade, sterile LB Agar plate. Seal in ziplock bag to keep bacteria and mold spores from air from the plates. This is especially important since antibiotic marker resistance is not being used. Leave out on bench top overnight (not placed in incubator because it isn't calibrated yet - this will be complete tomorrow). If the incubator is working and calibrated later this evening, they will be placed in the incubator.

8. Also started updating science lab inventory list. There are several things missing from the list that need to be included in the lab inventory.


EVA 10 Report

Christine Redmond Reporting

EVA Time: 11:55-14:05

Mode of Transportation: ATV

EVA Procedure: Crew members will travel on ATVs along Cow Dung Road, stopping about every two minutes for a radio signal strength check, recording GPS coordinates at each point of communication.

EVA Results: Traveled full length of trail in both the northern and southern directions and took radio signal measurements and GPS coordinates along full length of trail.

Lessons Learned: I need to wear the smaller helmet next time I go on a ATV EVA and remember to take earrings off. Also, for purposes similar to this EVA mission it may be helpful to count steps taken from road to get a more accurate distance from road approximation to report to base.



EVA 11 Report

Jon Kosh Reporting

EVA Crewmembers:

Mitesh Agrawal, KK4AQI

Jonathan Kosh, KF7GLD

Base Comms: Tommy Caillouet, KJ4KAC

EVA Time: 1643-1728

Mode of Transportation: Foot

EVA Procedure: Currently the radio repeater height is limited to ~10ft because the length of the coax cable is shorter than the length of the max fiberglass mast height. EVA 11's mission was to carry a coil of HF coax cable to the top of radio ridge and raise the repeater antenna to full height (25').  With high winds over the past week and increasingly higher winds expected the next, the team also carried a shovel to set the mast into the ground for more stability.

EVA Results: Operation not complete. Equipment successfully carried to the top of radio ridge, but in the process of breaking down the existing HF cable currently used to insert the longer coax cable, the repeater wires were removed from the back of their powerwerx connector. A new connector needed to be fit at the Hab. The team promptly returned to the Hab for a quick fix.

Lessons Learned: Make sure to rigorously test electrical connections created at the Hab before utilizing them in the field.



EVA 12 Report

Mitesh Agrawal Reporting

EVA Crewmembers:

Mitesh Agrawal, KK4AQI

Jonathan Kosh, KF7GLD

Base Comms: Tommy Caillouet, KJ4KAC

EVA Time: 1800-1915

Mode of Transportation: Foot

EVA Procedure: Team had to fix the loosened wire brought after EVA 11 operation, go to the top of radio ridge and raise the repeater antenna to the full height (25 feet)

EVA Results: Raised the height of the repeater antenna to the full height (25 feet)

Lessons Learned: Always better to take the tools required for some operation even if one suspects it might be slightly useful. We should have taken a hand drill to the ridge as initially, we could not fit our wires through one of the holes in the radio box but we managed to make it bigger using a screw driver.


Chef's Report

Christine Redmond Reporting

1a) Today was a non-cooking day

1b) Was today a special day celebrated at the main meal with special

food or activity?

No

2) List the foods served at today's main meal (usually the evening

meal), giving full names of each.

Santa Fe Black Beans and Rice

3) List any main meal foods not finished by the crew (leftovers stored for later use or discarded)

None

4) Recipes for anything more complicated than following package

directions exactly, or rehydrating.

5) Lessons learned (if any)

6) Comments/questions for Kim and Jean, the food study investigators


Journalist Report

Christine Redmond Reporting

We all live in a yellow submarine, a yellow submarine, a yellow submarine…Crew 101 is now half way through their mission! Has crew 101 gone crazy? You be the judge. Positive crew dynamics is essential for mission success. I am fortunate to be a part of a crew that works so well together, having many laughs along the way.

Today was a productive day on Mars despite windy conditions and losing power! This morning Cody and I headed out on EVA #10. We traveled the full length of the trail in both the northern and southern direction, stopping periodically for radio strength checks and GPS readings.  Cody and I had a blast navigating the rugged Martian terrain! One of the NASA INSPIRE students put together a study for the crew to evaluate the safety of the ATVs. Today was our first day participating in this study. Safety is top priority here in the isolated Maritain environment. Crew 101 always exercises appropriate safety measures.

After a quick lunch break, Mitesh and Jon were ready to suit up for EVA #11. This EVA involved climbing up radio ridge and raising the antenna for the repeater as well as further securing the antenna given the recent windy conditions. Jon and Mitesh were able to raise the antenna to 25 feet! We are all looking forward to testing our increased range during our EVAs tomorrow. We lost Hab power during this EVA. Tommy quickly moved to hand held radio communication and notified the EVA team of the outage.  Luckily, Cody (our chief engineer) sprung into action and discovered that power would be restored shortly.

Today is a pre-prepared day. Santa Fa Beans and Rice…let’s see how this one goes!